The number of customers that visit a local small business is 37,800 and has been continuously declining at a rate of 1.7% each year. What is the approximate number of customers that visit the business in 12 years
Answer by JBarnum(2044)

37800-(.017)(37800)=y1
37800-642.6=37157.4
37157.4-(.017)(37157.4)=y2
36525.7242-(.017)(36525.7242)=y3
35904.7869-(.017)(35904.7869)=y4
35294.4055-(.017)(35294.4055)=y5
34894.4006-(.017)(34894.4006)=y6
34101.1958
....
you could keep going to 12
i went blank sry i cant remember the formula to find this answer faster